Suppose that in July​ 2013, Nike Inc. had sales of

$25,339 million, EBITDA of $3,258 ​million, excess cash of $3,335​million, $1,383 million of​ debt, and 884.9 million shares outstanding.

P/E

Price / Book

Enterprise Value / Sales

Enterprise Value / EBITDA

Average

29.84

2.44

1.12

9.76

Maximum

+136​%

+70​%

+55​%

+86​%

Minimum

-62​%

−63​%

-48​%

-34​%

a. Using the average enterprise value to sales multiple in the table​ above, estimate​ Nike's share price.

b. What range of share prices do you estimate based on the highest and lowest enterprise value to sales multiples in the table​ above?

c. Using the average enterprise value to EBITDA multiple in the table​ above, estimate​ Nike's share price.

d. What range of share prices do you estimate based on the highest and lowest enterprise value to EBITDA multiples in the table​ above?
Answer #1

1.
=(1.12*25339+3335-1383)/884.9=34.2769578483445

2.
Lowest=(1.12*(1-48%)*25339+3335-1383)/884.9=18.882849587524

3.
Highest=(1.12*(1+55%)*25339+3335-1383)/884.9=51.9160402305345

4.
=(9.76*3258+3335-1383)/884.9=38.1399932195728

5.
Lowest=(9.76*(1-34%)*3258+3335-1383)/884.9=25.922401175274

6.
Highest=(9.76*(1+86%)*3258+3335-1383)/884.9=69.0433142727992